USING the uncertainty principle,
estimate the minimum energy of a particle in a simple harmonic potential U= ½kx² .
1
Expert's answer
2020-01-06T10:16:36-0500

The energy value can be calculated on the basis of the Hamiltonian of the system as follows:

"E = <\hat{H}>=<\hat{T}> + <\hat{U}> = \frac{<\hat{p}^2>}{2m} + \frac{k <\hat{x}^2>}{2}"

By definition, the uncertainties of the momentum and position operators are equal to:

"(\Delta{p})^2 = <\hat{p}^2> - <\hat{p}>^2, \quad (\Delta{x})^2 = <\hat{x}^2> - <\hat{x}>^2"

Due to the symmetry of potential "<\hat{p}> = 0 = <\hat{x}>" .

Hence, we obtain:

"E = \frac{(\Delta p)^2}{2m} + \frac{k (\Delta x)^2}{2}" .

The uncertainty principle states the following: "\Delta p \Delta x \geq \frac{\hbar}{2}" . Hence, "\Delta p \geq \frac{\hbar}{2 \Delta x}" . As a result:

"E \geq \frac{(\hbar)^2}{8 m (\Delta x)^2} + \frac{k (\Delta x)^2}{2}"

The expression on the right side is minimal when "\frac{d}{dx} \left[ \frac{(\hbar)^2}{8 m (\Delta x)^2} + \frac{k (\Delta x)^2}{2} \right] = 0". Doing this explicitly leads to the following result: "(\Delta x)^2 = \frac{\hbar}{2 \sqrt{km}}=\frac{\hbar}{2 m \omega}," where we utilized the relation "\omega^2 = \frac{k}{m}." Substituting this into the expression for energy results in the following:

"E_{min}= \frac{\hbar^2 2 m\omega}{8m \hbar}+ \frac{m \omega^2 \hbar}{4 m \omega} = \frac{\hbar \omega}{2}"